OpenStudy (anonymous):

For what positive integers k is the following series convergent? \[\sum_{n=1}^{\infty}{\frac{(n!)^2}{(kn)!}}\]

OpenStudy (zarkon):

looks like \[k\ge2\]

OpenStudy (anonymous):

yes using the ratio test.

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

how do you "simplify" \[\frac{(kn)!}{(kn+k)!}\]? i know this goes to zero very rapidly, but can't figure out a nice form for it

OpenStudy (zarkon):

\[\frac{1}{(kn+1)(kn+2)\cdots (kn+k)}\]

OpenStudy (anonymous):

just 'stretch' out the factorial so it cancels

OpenStudy (anonymous):

so you get \[\frac{(n+1)^2}{(kn+1)(kn+2)\cdots (kn+k)}\] by ratio test or did i make a mistake?

OpenStudy (zarkon):

that is correct